A 2.5 gram penny is given a charge of -4.0 *10^-9 c. How many excess electrons are on the penny?
Answer #1

Using the quantisation of charge ,Q=ne

4*10^-9=n*1.6*10^-19

n=2.5*10^10 is no. of excess electrons